# Maximum Entropy Temporal Networks

### v4

@@ Abstract
These datasets span domains from organizational communication to international relations, physical proximity, and online social media. They also differ in density, timescale, and heterogeneity. The common challenge is to separate the features of each dataset—such as degree heterogeneity or daily cycles—from significant temporal signatures, such as anomalous motifs, excess reciprocity, or collective bursts, by defining ensembles that enforce the known marginals (e.g., strengths, block structure) and temporal baselines (Poisson or Hawkes profiles) while leaving higher-order correlations unconstrained. By generating expectations and variances for the relevant statistics under these ensembles, we can define the associated empirical deviations.
In the Enron dataset we test that the frequency of reciprocated motifs largely exceeds the maximum-entropy expectation given node strengths and bursty Hawkes timing. In the MID data, one can assess whether conflict proliferation across specific edges or blocks is more common than expected given the observed distribution of disputes. In Reality Mining, deviations from the null can highlight the role of small-group gatherings and scheduled rhythms, while in Facebook the null benchmark reveals when extreme broadcast behavior of hubs is attributable to degree heterogeneity alone or signals additional bursty mechanisms at a node level. In each case, the ensemble serves as the baseline against which significance is judged: if an observed property falls within the distribution generated by the maximum-entropy and NHPP model, it is explained by the enforced constraints; if not, it signals mechanisms beyond those constraints.
Figure 1: Time-dependent motif ratio for reciprocated links in the Enron dataset versus the maximum-entropy temporal network ensembles obtained combining Poisson/Exp-Hawkes (GH-Exp)/PL-Hawkes (GH-PL) temporal profiles and configuration model (CM) and block-configuration model (Block-CM) edge structures.
Discussion We have introduced a maximum-entropy framework for temporal networks in continuous time, combining inhomogeneous Poisson processes with static maximum-entropy edge distributions. The key consequence of the derivation is a clean factorization of time and marks, which yields closed-form likelihoods, tractable fitting procedures, and explicit formulas for edge coverage, burstiness, and motif statistics. The framework provides a benchmark for assessing which features are explained by basic structural and temporal constraints and which require additional mechanisms. In particular, the use of NHPP improves agreement with empirical inter-event distributions and motif counts, while the static maximum-entropy edge modeling accounts for degree and block structure. The resulting ensembles thus serve as principled null models for temporal networks, offering a consistent basis for significance testing and model comparison. The present results allow immediate generalizations in several directions. The use of inhomogeneous Poisson processes enables us to fit an observed bursty behavior in the data or to impose it altogether with a theoretical model. In the maximum-entropy construction here presented there is no direct self-exciting behavior in the dataset generation, i.e. the self-excitation is driven by the original timestamped process. One direction is to explore richer time profiles, including explicit calibration with renewal processes, block-specific multivariate Hawkes processes, and neural kernel estimation calibration, bridging the gap with graph-neural network approaches. The framework can be used to evaluate higher-order structures beyond two-event motifs, providing a systematic baseline for studying temporal clustering, cascades, or community dynamics. In all these cases, the maximum-entropy perspective clarifies what can be expected from constraints alone and where genuinely new patterns emerge in networked systems changing continuously over time.
Supplementary Material: Maximum Entropy Temporal Networks
I Marked point processes, likelihood, and MaxEnt
Events and marks.
Let $\mathcal{V}=\{1,\dots,n\}$ and $E\subseteq\{(i,j)\in\mathcal{V}^{2}:i\neq j\}$ be the set of admissible directed edges. A temporal network is a marked point process on $[0,T]$ with events $\{(t_{k},m_{k})\}_{k=1}^{K}$, where $t_{k}\in(0,T]$ and $m_{k}=(i_{k}\!\to\!j_{k})\in E$. We denote by $N_{ij}(I)$ the count of $(i\!\to\!j)$ events whose times lie in an interval $I\subseteq[0,T]$, and by $K(I)=\sum_{(i,j)\in E}N_{ij}(I)$ the total event count.
Factorized intensity (time $\times$ marks).
We fix a partition $P=\{E_{r}\}_{r=1}^{R}$ of $E$. The model intensity factorizes as
$\lambda_{ij}(t)=\phi_{r(i,j)}(t)\,w_{ij},\qquad(i,j)\in E,$ (12)
where: (i) $\phi_{r}(t)\geq 0$ is the time layer for part $r$; (ii) $w_{ij}\geq 0$ are time–independent mark weights (with $w_{ii}=0$). Define the per–part mass and integral
$W_{r}:=\sum_{(i,j)\in E_{r}}w_{ij},\qquad\Lambda_{r}(I):=\int_{I}\phi_{r}(t)\,W_{r}\,\mathrm{d}t,$ (13)
and the normalized mark law in part $r$,
$\Pi^{(r)}_{ij}:=\begin{cases}\dfrac{w_{ij}}{W_{r}},&(i,j)\in E_{r},\\[6.0pt] 0,&\text{otherwise}.\end{cases}\qquad\sum_{(i,j)}\Pi^{(r)}_{ij}=1.$ (14)
Then (12) can be re-written as:
$\lambda_{ij}(t)=\Pi^{(r)}_{ij}\,\lambda_{r}(t),\qquad\lambda_{s}(t):=\phi_{r}(t)\,W_{r}.$ (15)
Likelihood and exact split.
For a marked nonhomogeneous Poisson process, the log–likelihood is
$\ell(\{\lambda_{ij}\})=\sum_{k=1}^{K}\log\lambda_{i_{k}j_{k}}(t_{k})-\sum_{(i,j)\in E}\int_{0}^{T}\lambda_{ij}(t)\,\mathrm{d}t.$ (16)
If (12) holds globally ($R=1$, write $\phi,\;w$), then
$\ell(\phi,w)=\sum_{k=1}^{K}\log\phi(t_{k})-W\!\int_{0}^{T}\phi(t)\,\mathrm{d}t+\sum_{(i,j)}N_{ij}(I)\log w_{ij},$ (17)
which separates time (first two terms) and marks (last term). For a general partition, the split reads
$\ell(\{\phi_{r}\},w)=\sum_{r=1}^{R}\Big(\sum_{k:\,m_{k}\in E_{r}}\log\phi_{r}(t_{k})-W_{r}\!\int_{0}^{T}\phi_{r}(t)\,\mathrm{d}t\Big)+\sum_{(i,j)}N_{ij}(I)\log w_{ij}.$ (18)
Maximum-entropy derivation and factorization theorem.
Maximizing path entropy (equivalently, minimizing $\sum_{ij}\int(\lambda_{ij}\log\lambda_{ij}-\lambda_{ij})dt$) subject to linear constraints that (i) fix partition totals in time,
$\sum_{(i,j)\in E_{r}}\lambda_{ij}(t)=G_{r}(t)\quad\forall r,\;t,$ (19)
and (ii) fix time–integrated margins (expected edge totals $\mu_{ij}$, node strengths $s_{i}^{\text{out}}=\sum_{j}N_{ij}$, $s_{j}^{\text{in}}=\sum_{i}N_{ij}$, block totals $N_{ab}$, etc.), yields the Euler–Lagrange stationarity condition that leads to the factorization
$\lambda_{ij}(t)=\exp\!\Big\{\alpha_{r(i,j)}(t)+\Theta_{ij}+\theta_{i}^{\rm out}+\theta_{j}^{\rm in}+\zeta_{g(i),g(j)}-1\Big\}\!,$ (20)
i.e. the product form (12) with $\phi_{r}(t)=e^{\alpha_{r}(t)}$ and $w_{ij}=\exp(\Theta_{ij}+\theta_{i}^{\rm out}+\theta_{j}^{\rm in}+\zeta_{g(i),g(j)}-1)$. Thus the factorization is the unique maximum-entropy solution for linear time and mark constraints. [^1] Forbidden edges can be enforced by a mask $M_{ij}\in\{0,1\}$ by adding $\log M_{ij}$ to the mark potential.
I.1 Fitted marks layer
Given $G_{r}$ (or a parametric fit for $\lambda_{r}$) and chosen margins, one solves for $\Pi$ via masked biproportional scaling (IPFP/Sinkhorn):
Edge totals: $\int_{0}^{T}\lambda_{ij}(t)\,\mathrm{d}t=\mu_{ij}\;\;\Rightarrow\;\;\Pi_{ij}=\mu_{ij}/\Lambda^{tot}.$ (21)
Strengths: $\sum_{j\neq i}\Pi_{ij}=\frac{s_{i}^{\rm out}}{\Lambda^{tot}},\qquad\sum_{i\neq j}\Pi_{ij}=\frac{s_{j}^{\rm in}}{\Lambda^{tot}},$ (22)
where we are allocating the events generated by the time profiles $\lambda_{r}(t)$ across the available edges according to the matrix $\Pi$.
I.2 Fitted time layers (Poisson/Hawkes)
For stationary linear Hawkes $\lambda_{r}(t)=\mu_{r}+\int_{0}^{\infty}g_{r}(u)\,\mathrm{d}N_{r}(t-u)$ with branching ratio $\eta_{r}:=\int_{0}^{\infty}g_{r}(u)\mathrm{d}u<1$, one has the stationary identities, [24, 25, 26],
$\bar{\lambda}_{r}=\frac{\mu_{r}}{1-\eta_{r}},\qquad S_{r}(0)=\frac{\bar{\lambda}_{r}}{(1-\eta_{r})^{2}},\qquad\bar{\lambda}_{r}+2\!\int_{0}^{\infty}\gamma_{r}(u)\,\mathrm{d}u=S_{r}(0),$ (23)
where we introduced the Bartlett spectrum, $S_{r}(\omega)=\bar{\lambda}_{r}\big|1-\mathcal{G}_{r}(\omega)\big|^{-2}$, with $\mathcal{G}_{r}(\omega)=\int_{0}^{\infty}g_{r}(u)e^{-i\omega u}\mathrm{d}u$ and for stationary linear Hawkes models, $\gamma_{r}$ is the inverse Fourier transform of the Bartlett spectrum. Main linear Hawkes cases:
Poisson: $\quad g_{r}\equiv 0\Rightarrow\eta_{r}=0,\;\bar{\lambda}_{r}=\mu_{r},\;S_{r}(0)=\bar{\lambda}_{r},\;\gamma_{r}\equiv 0.$ (24)
[…]
For a kernel $g_{r}(u)=\alpha_{r}e^{-\beta_{r}u}$ with branching ratio $n_{r}=\alpha_{r}/\beta_{r}<1$ and stationary rate $\bar{\lambda}_{r}=\mu_{r}/(1-n_{r})$, the intensity autocovariance decays exponentially,
$\gamma_{r}(u)=\bar{\lambda}_{r}\,\frac{\alpha_{r}^{2}}{\beta_{r}-\alpha_{r}}\,e^{-(\beta_{r}-\alpha_{r})|u|}.$ (27)
In the maximum-entropy construction, once we fit the arbitrary point processes to the observed series of timestamps we use that as a constraint for the inhomogeneous Poisson process, i.e. the timestamp generation is then history-independent in the maximum-entropy case. Similarly, one can reproduce the renewal density of a renewal process with a history-independent inhomogeneous Poisson process.
II Partition models
To combine structural and temporal heterogeneity, we couple distinct mark models (encoding edge–weight constraints) with different temporal kernels. Each partition type uses a consistent mark allocation but allows for different time dynamics. Table 1 summarizes all model combinations.
Table 1: Combinations of mark and time models across partition types.
Partition Mark (edge–weight) model Time model
Global WCM (Weighted configuration model) Poisson / Hawkes–Exp / Hawkes–PL
Block–pair Blockwise WCM (block $\times$ block expected weights) Poisson / Hawkes–Exp / Hawkes–PL
Static baseline WCM Homogeneous Poisson (no excitation)
Block–pair partition.
Nodes are first assigned to structural communities or blocks (e.g. via Louvain modularity or metadata labels). Each ordered block pair $(a,b)$ forms a subpartition $r=(a,b)$ with its own temporal kernel $g_{r}(u)$ and mark matrix $\Pi^{(r)}_{ij}$ constrained by blockwise-WCM expectations. This representation preserves block–to–block strength constraints and allows distinct excitation dynamics across inter– and intra–block activity. Block pairs therefore capture mesoscale heterogeneity that is missed by global or node–level partitions.
II.1 Frozen–path NHPP surrogates
In the maximum-entropy construction, we are maximizing the entropy of a non-homogeneous Poisson process. Nevertheless in many practical cases, we would like to model the train of events with history-dependent stochastic processes, such as Hawkes process, doubly stochastic Poisson processes (Cox process), or renewal processes. Self–exciting processes such as the linear Hawkes model with exponential kernel,
$\lambda(t)=\mu+\sum_{t_{k}<t}\alpha\,e^{-\beta(t-t_{k})},\qquad\eta=\alpha/\beta<1,$ (28)
introduce stochastic feedback: each event transiently raises the likelihood of future events. The branching ratio $\eta$ controls the expected number of triggered descendants per event, and $\beta^{-1}$ sets the memory decay time. In order to maintain the relevant properties of the stochastic models we can introduce frozen–path NHPP surrogates, which preserve the fitted time–dependent rate $\lambda(t)$ of a stochastic model (e.g., Hawkes or renewal process) but remove stochastic feedback. Formally, one replaces the random history-dependent intensity $\lambda(t|\mathcal{H}_{t})$ by a deterministic path $\hat{\lambda}(t)=\lambda(t|\mathcal{H}_{t})$ and re-samples events as an NHPP with this frozen rate. This surrogate maintains the expected counts $\Lambda(t)$, diurnal or seasonal modulations, and mark allocation structure, but suppresses the explicit stochastic feedback between events. In general, for the individual frozen process for a given interval,
$\int_{a}^{b}\!\!\int_{0}^{\Delta}\hat{\lambda}(t)\hat{\lambda}(t+u)\,du\,dt\neq\int_{a}^{b}\!\!\int_{0}^{\Delta}\mathbb{E}[\lambda(t)\lambda(t+u)]\,du\,dt.$ (29)
However, when averaged over realizations of $\lambda(t)$ drawn from the original process, the expected pair–correlation of frozen NHPPs equals that of the full process:
$\mathbb{E}_{\lambda}\!\left[\int_{a}^{b}\!\!\int_{0}^{\Delta}\hat{\lambda}(t)\hat{\lambda}(t+u)\,du\,dt\right]\sim\int_{a}^{b}\!\!\int_{0}^{\Delta}\mathbb{E}[\lambda(t)\lambda(t+u)]\,du\,dt,$ (30)
i.e. when averaging over realizations of the original point process the frozen autocovariance approximates the theoretical stationary autocovariance of the stochastic point process. Thus, the frozen–path NHPP preserves the mean intensity and motif time factors in expectation, though any given realization lacks endogenous stochastic clustering.
Figure 2: Frozen-path NHPP runs approximating the theoretical Hawkes auto-covariance.
In renewal models, the inter–event distribution itself is specified rather than generated by feedback. Hazard rates depend on the elapsed time since the previous event and can reproduce over–dispersed inter-event times. Similarly, we can convert a fitted renewal process into a frozen NHPP surrogate removing the renewal memory yet keeping the original inter-event times on average.
II.2 Inter–event times, rate variability, and burstiness
For each edge $(i,j)$, let $\Delta_{ij}$ denote the inter–event time, or inter–spike interval (ISI), i.e., the time difference between two consecutive events on that edge. Inter–event–time distributions summarize the microscopic temporal organization of event sequences and connect directly to notions of regularity and burstiness.
For a stationary Poisson process with constant rate $\bar{\lambda}_{ij}$, events are independent and identically distributed, and inter–event times are exponentially distributed:
$p(\Delta_{ij})=\bar{\lambda}_{ij}\,e^{-\bar{\lambda}_{ij}\Delta_{ij}},\qquad\mathbb{E}[\Delta_{ij}]=\frac{1}{\bar{\lambda}_{ij}},\qquad\sigma^{2}[\Delta_{ij}]=\frac{1}{\bar{\lambda}_{ij}^{2}}.$ (31)
The process is memoryless, with coefficient of variation (CV) equal to one. To quantify deviations from this baseline, define the coefficient of variation and the burstiness index:
$\mathrm{CV}_{ij}=\frac{\sqrt{\sigma^{2}[\Delta_{ij}]}}{\mathbb{E}[\Delta_{ij}]},\qquad B_{ij}=\frac{\mathrm{CV}_{ij}-1}{\mathrm{CV}_{ij}+1},$ (32)
where $B_{ij}=0$ for a Poisson process, $B_{ij}>0$ for bursty dynamics (over–dispersion), and $B_{ij}<0$ for regular, quasi–periodic dynamics.
If the instantaneous rate $\lambda_{ij}(t)$ varies over time, events remain conditionally independent given $\lambda_{ij}(t)$, but the process is no longer stationary. Such a process is a nonhomogeneous Poisson process (NHPP), defined by its deterministic intensity $\lambda_{ij}(t)\geq 0$ and compensator $\Lambda_{ij}(t)=\int_{0}^{t}\lambda_{ij}(s)\,ds$. Periods of high activity can produce clusters of short intervals, while low–rate periods generate long gaps. Hence, a NHPP can exhibit apparent burstiness ($B_{ij}>0$) arising entirely from deterministic rate modulation, even though events are conditionally Poisson and possess no genuine memory. Thus, renewal and Hawkes models can generate identical mean rates but differ sharply in their conditional ISI structure. Burstiness in event timing can therefore originate from two fundamentally distinct mechanisms:
  • Exogenous modulation—deterministic variation in $\lambda(t)$ (as in a NHPP or frozen–path surrogate), which produces clustered intervals via rate heterogeneity but no true conditional dependence; and
  • Endogenous excitation—stochastic feedback (as in Hawkes or renewal dynamics), which generates genuine temporal correlations and amplifies variance even at fixed rate.
Both mechanisms yield broad inter–event–time distributions and $B_{ij}>0$, but only the latter reflects intrinsic memory in the event sequence. The comparative analysis of homogeneous Poisson, time–varying NHPP, renewal, and Hawkes models thus disentangles deterministic rate fluctuations from stochastic self–excitation as sources of bursty dynamics.
III Explicit expectations in time $\times$ marks form
All expectations below depend only on $(\lambda_{r})$ and $(\Pi^{(r)})$ via (15).
III.1 Counts, strengths, rates
For any interval $I=[a,b]\subseteq[0,T]$,
$\mathbb{E}\big[N_{ij}(I)\big]$ $=\Pi^{(r)}_{ij}\;\Lambda_{r}(I),\,\,(i,j)\in E_{r}$ (33)
$\mathbb{E}\big[K(I)\big]$ $=\sum_{(i,j)}\mathbb{E}[N_{ij}(I)]=\sum_{r=1}^{R}\Lambda_{r}(I),$ (34)
$\mathbb{E}\big[s_{i}^{\rm out}(I)\big]$ $=\sum_{j}\mathbb{E}[N_{ij}(I)]=\sum_{s}\Lambda_{s}(I)\delta_{r(i,j),s}\sum_{j}\Pi^{(s)}_{ij}$, (35)
$\mathbb{E}\big[s_{j}^{\rm in}(I)\big]$ $=\sum_{i}\mathbb{E}[N_{ij}(I)]=\sum_{r}\Lambda_{r}(I)\sum_{i}\Pi^{(r)}_{ij}$. (36)
If each $\lambda_{r}$ is (wide–sense) stationary with mean $\bar{\lambda}_{r}$, the long–run edge rate is
$\bar{\lambda}_{ij}:=\lim_{|I|\to\infty}\frac{\mathbb{E}[N_{ij}(I)]}{|I|}=\Pi^{(r)}_{ij}\,\bar{\lambda}_{r}.$ (37)
III.2 Expected unique directed edges in II
Under conditional Poisson arrivals, the per–edge count over $I$ is Poisson with mean $\mu_{ij}(I):=\mathbb{E}[N_{ij}(I)]$ from (33). Thus the expected number of distinct directed edges observed in $I$ is
$\mathbb{E}[U_{\to}(I)]=\sum_{i\neq j}\big(1-e^{-\mu_{ij}(I)}\big),\qquad\mu_{ij}(I)=\Pi^{(r)}_{ij}\,\Lambda_{r}(I).$ (38)
Block–pair unique–edge expectations follow by summation over $(i,j)$ with $g(i)=a,g(j)=b$.
III.3 Motif pairs in a tolerance window $\Delta$
Consider ordered pairs of events with lag $u\in(0,\Delta]$. For two parts $r,s$ define sender/receiver masses
$r^{(r)}_{i}:=\sum_{j}\Pi^{(r)}_{ij},\qquad c^{(r)}_{j}:=\sum_{i}\Pi^{(r)}_{ij}$. (39)
The pure–mark probabilities, i.e. the chance that the second edge satisfies a motif given the first edge, are
$p_{\textsc{rep}}^{(r,s)}$ $=\sum_{i,j}\Pi^{(r)}_{ij}\,\Pi^{(s)}_{ij}$, $p_{\textsc{rec}}^{(r,s)}$ $=\sum_{i,j}\Pi^{(r)}_{ij}\,\Pi^{(s)}_{ji}$, (40)
$p_{\textsc{bcast}}^{(r,s)}$ $=\sum_{i}r^{(r)}_{i}\,r^{(s)}_{i}-p_{\textsc{rep}}^{(r,s)},$ $p_{\textsc{conv}}^{(r,s)}$ $=\sum_{j}c^{(r)}_{j}\,c^{(s)}_{j}-p_{\textsc{rep}}^{(r,s)}.$ (41)
The time contribution is the per–part intensity cross–moment integrated over the window:
$C_{r,s}(\Delta;I):=\int_{a}^{b}\int_{0}^{\Delta}\mathbb{E}\big[\lambda_{r}(t)\lambda_{s}(t+u)\big]\,\mathrm{d}u\,\mathrm{d}t,\qquad I=[a,b].$ (42)
Then the expected motif counts are
$\boxed{\;\mathbb{E}\big[M_{\mathrm{TYPE}}^{(\Delta)}(I)\big]=\sum_{r,s=1}^{R}p_{\mathrm{TYPE}}^{(r,s)}\,C_{r,s}(\Delta;I),\quad\mathrm{TYPE}\in\{\textsc{rec},\textsc{bcast},\textsc{conv},\textsc{rep}\}.\;}$ (43)
Two explicit evaluations are useful:
  • If $r\neq s$ are independent layers, $\mathbb{E}[\lambda_{r}(t)\lambda_{s}(t+u)]=\bar{\lambda}_{r}\,\bar{\lambda}_{s}$ and
    $C_{r,s}(\Delta;I)=|I|\,\Delta\,\bar{\lambda}_{r}\,\bar{\lambda}_{s}.$ (44)
  • For $r=s$, decompose $\mathbb{E}[\lambda_{r}(t)\lambda_{r}(t+u)]=\bar{\lambda}_{r}^{2}+\gamma_{r}(u)$ to get
    $C_{r,r}(\Delta;I)=|I|\Big(\bar{\lambda}_{r}^{2}\,\Delta+\int_{0}^{\Delta}\gamma_{r}(u)\,\mathrm{d}u\Big).$ (45)
Integrating (27) in (45) gives
$C_{r,r}(\Delta;I)=|I|\!\left[\bar{\lambda}_{r}^{2}\,\Delta+\bar{\lambda}_{r}\,\frac{\alpha_{r}^{2}}{(\beta_{r}-\alpha_{r})^{2}}\big(1-e^{-(\beta_{r}-\alpha_{r})\Delta}\big)\right].$ (46)
The first term represents the Poisson baseline, while the second quantifies the temporal amplification due to self–excitation, which grows with the excitation strength $\alpha_{r}$ and the kernel memory $\beta_{r}^{-1}$.
III.4 Model selection checks and identifiability
Scaling invariance.
If $w\mapsto c\,w$ and $\phi_{r}\mapsto\phi_{r}/c$ for all $r$ (or $W_{r}\mapsto cW_{r}$ and $\phi_{r}\mapsto\phi_{r}/c$), then $\lambda_{ij}$ is unchanged. Any fitting scheme must fix this degree of freedom, e.g. by working with $(\lambda_{r},\Pi^{(r)})$ directly, or by fixing $W_{r}$ (or $\int\phi_{r}$).
Feasibility of margins.
Strength/block totals must be compatible with the mask and per–part masses. A necessary condition is that each part saturates: $\sum_{(i,j)\in E_{r}}\mu_{ij}^{\star}=\Lambda_{r}([0,T])$; row/column sums must agree across decompositions. IPFP converges iff the mask admits a feasible point with those margins.
Stability of Hawkes layers.
Each Hawkes layer must satisfy $\eta_{r}<1$ to have finite $\bar{\lambda}_{r}$ and $S_{r}(0)$; otherwise expectations in §II.2–§III.3 are ill–posed.
III.5 Structural and temporal analyses (Enron)
Figure 3: Degree-related statistics for Enron TRAIN (blockpair–PL model). observed vs. expected out-degree
Figure 4: Burstiness vs. degree for the Enron TRAIN dataset. The plot compares empirical observations with the blockpair–PL model, showing how local transitivity scales with node degree.
Figure 5: Log of event counts per edge (left) and Burstiness per edge (right) for the Enron TRAIN dataset.
Figure 6: Block-to-block edge count comparison for Enron TRAIN: (left) empirical unique edges, (right) expected under the Block–DC–DWCM constraint.
Figure 7: Inter-event time distributions (linear scale) for the Enron TRAIN dataset, comparing empirical data with fitted temporal models. The overlay highlights how each process reproduces or underestimates temporal burstiness.
Figure 8: Raster plots of the top-5 intra-block event sequences for Enron TRAIN: (left) empirical; (right) blockpair–PL model (sample s=00s{=}00).
Figure 9: Enron TRAIN lambda over time Poisson CM ME
Figure 10: Enron TRAIN lambda over time GH
@@ III.5 Structural and temporal analyses (Enron)
Figure 11: Enron TRAIN motif ratios (mean $\pm$ s.d.) over time for $\Delta_{\text{large}}$.
(a) CON, $\Delta_{\text{large}}$
(b) REC, $\Delta_{\text{large}}$
(c) REP, $\Delta_{\text{large}}$
(d) BRO, $\Delta_{\text{large}}$
Figure 12: Reality mining TRAIN motif ratios (mean $\pm$ s.d.) over time for $\Delta_{\text{large}}$.
(a) CON, $\Delta_{\text{large}}$
(b) REC, $\Delta_{\text{large}}$
(c) REP, $\Delta_{\text{large}}$
(d) BRO, $\Delta_{\text{large}}$
Figure 13: Militarized Interstate Disputes TRAIN motif ratios (mean $\pm$ s.d.) over time for $\Delta_{\text{large}}$.
(a) CON, $\Delta_{\text{large}}$
(b) REC, $\Delta_{\text{large}}$
(c) REP, $\Delta_{\text{large}}$
(d) BRO, $\Delta_{\text{large}}$
Figure 14: Facebook TRAIN motif ratios (mean $\pm$ s.d.) over time for $\Delta_{\text{large}}$.
Table 2: Per-event log-likelihoods for the Enron dataset under different temporal models. LLmark/ev refers to the mark (dyad) likelihood, LLtime/ev to the temporal likelihood, and LLtotal/ev to their sum.
Partition Time model LLmark/ev LLtime/ev LLtotal/ev
Global Poisson 7.53-7.53 $0.283$ 7.25-7.25
Global Hawkes-Exp 7.53-7.53 $0.291$ 7.24-7.24
Global Hawkes-PL 7.53-7.53 $0.775$ 6.76-6.76
Blockpair Poisson 3.45-3.45 2.10-2.10 5.55-5.55
Blockpair Hawkes-Exp 3.45-3.45 2.06-2.06 5.50-5.50
Blockpair Hawkes-PL 3.45-3.45 1.61-1.61 5.06-5.06
Table 3: Comparison of empirical and model-generated temporal-network statistics. Values are shown as mean $\pm$ standard deviation across Monte Carlo samples.
Model Nodes Events Unique Edges ISI Mean CV Burstiness
Empirical 136 2999 513 0.277 2.74 0.465
Poisson+CM $135.5\pm 0.5$ $2990\pm 60$ $1450\pm 20$ $0.279\pm 0.005$ $991\pm 0.016 0.0$$004\pm 0.008$
GH-Exp+CM $135.3\pm 0.5$ $2700\pm 60$ $1350\pm 30$ $0.308\pm 0.007$ $1.00\pm 0.02$ $0.001\pm 0.011$
[…]
GH-PL+CM $135.4\pm 0.5$ $2700\pm 40$ $1349\pm 18$ $0.309\pm 0.005$ $1.23\pm 0.02$ $0.105\pm 0.009$
GH-PL+Block-CM $135.6\pm 0.5$ $2510\pm 50$ $697\pm 11$ $0.331\pm 0.007$ $1.04\pm 0.03$ $0.020\pm 0.013$
Table 4: Enron TRAIN: model summary. For each model, the table reports the number of nodes, total number of events, number of unique directed edges, mean inter-event time (ISI_mean), coefficient of variation (CV), and burstiness index. Values are shown as mean ± standard deviation across generated samples.
Table 5: Comparison of empirical and model-generated motif statistics. Values represent mean $\pm$ standard deviation across Monte Carlo samples.
Model REP REC CON BRO
Empirical 1.17 0.40 0.29 0.624
Poisson+CM $034\pm 0.004 0.0$$010\pm 0.002$ $0.56\pm 0.02$ $0.77\pm 0.04$
GH-Exp+CM $031\pm 0.004 0.0$$0096\pm 0.0020$ $0.51\pm 0.03$ $0.71\pm 0.03$
GH-Exp+Block-CM $0.41\pm 0.03$ $0.12\pm 0.02$ $0.23\pm 0.02$ $0.34\pm 0.02$
GH-PL+CM $033\pm 0.003 0.0$$0105\pm 0.0019$ $0.552\pm 0.020$ $0.76\pm 0.03$
GH-PL+Block-CM $185\pm 0.015 0.0$$075\pm 0.006$ $0.217\pm 0.007$ $0.34\pm 0.02$
Table 6: Temporal motif ratios (Enron TRAIN, $\Delta_{\text{large}}$). Ratios are computed as the normalized counts of reciprocal (REP), repeated (REC), convergent (CON), and broadcast (BRO) motifs within the large temporal window $\Delta_{\text{large}}$.

### v5

@@ III Results
The challenge is to separate the features of the dataset —such as degree heterogeneity or daily cycles— from significant temporal signatures, such as anomalous motifs, excess reciprocity, or collective bursts, by defining ensembles that enforce the known marginals (e.g., strengths, block structure) and temporal baselines (Poisson or Hawkes profiles) while leaving higher-order correlations unconstrained. By generating expectations and variances for the relevant statistics under these ensembles, we can define the associated empirical deviations.
We test that the frequency of reciprocated motifs largely exceeds the maximum-entropy expectation given node strengths and bursty Hawkes timing. This can be simply interpreted as people working at Enron replying to each other more than would be expected simply from being active and emailing in bursts, thus suggesting genuine conversational feedback between workers. Figure 4 reports the time evolution of the four directed motifs under $\Delta_{\text{large}}$ for the Enron TRAIN split, comparing empirical ratios with the maximum-entropy temporal ensembles. Reciprocity (REC) and repetition (REP) are substantially underestimated by configuration-only models, even when coupled with Hawkes timing, indicating that strength constraints and global burstiness alone are insufficient to reproduce dyadic memory effects. Introducing block structure partially corrects this gap at short times but still decays away from the empirical curve, suggesting that mesoscale structure explains part—but not all—of the excess reciprocity. For convergence (CON) and broadcast (BRO), Hawkes-based models improve the early-time amplification relative to Poisson baselines, yet persistent discrepancies remain, especially at longer horizons where empirical ratios stabilise above the ensemble means. Overall, the figure highlights that while the time–mark factorization captures bursty temporal activity and degree constraints, motif persistence in the data reflects additional temporal correlations beyond those encoded by global Hawkes excitation and static block masks. The ensemble serves as the baseline against which significance is judged: if an observed property falls within the distribution generated by the maximum-entropy temporal model, it is explained by the enforced constraints; if not, it signals mechanisms beyond those constraints.
@@
Figure 4: Time-dependent motif ratios for the Enron dataset versus the maximum-entropy temporal network ensembles obtained combining Poisson/Exp-Hawkes (GH-Exp)/PL-Hawkes (GH-PL) temporal profiles and configuration model (CM) and block-configuration model (Block-CM) edge structures.
IV Conclusion
We have introduced a maximum-entropy framework for temporal networks in continuous time, combining non-homogeneous Poisson processes with static maximum-entropy edge distributions. The key consequence of the derivation is a clean factorization of time and marks, which yields closed-form likelihoods, tractable fitting procedures, and explicit formulas for edge coverage, burstiness, and motif statistics. The framework provides a benchmark for assessing which features are explained by basic structural and temporal constraints and which require additional mechanisms. In particular, the use of NHPP improves agreement with empirical inter-event distributions and motif counts, while the static maximum-entropy edge modeling accounts for degree and block structure. The resulting ensembles thus serve as principled null models for temporal networks, offering a consistent basis for significance testing and model comparison. The present results allow immediate generalizations in several directions. The use of non-homogeneous Poisson processes enables us to fit an observed bursty behavior in the data or to impose it altogether with a theoretical model. In the maximum-entropy construction here presented there is no direct self-exciting behavior in the dataset generation, i.e. the self-excitation is driven by the original timestamped process. The introduction of this ensemble connecting temporal networks to non-homogeneous Poisson processes provides a framework to naturally explore richer time profiles in networks: accounting for burstiness-degree correlations by introducing network-dependent factors in the intensity rates, including explicit heterogeneous calibration of the inter-event time distribution based on renewal processes theory, fitting block- and node-specific multivariate Hawkes processes, and generalising the kernel estimation to neural network functions [17], thus contributing to bridging the gap between maximum entropy network ensembles and temporal graph-neural networks [24]. The framework can be used to evaluate higher-order structures beyond two-event motifs, providing a systematic baseline for studying temporal clustering, cascades, or community dynamics. In all these cases, the maximum-entropy perspective clarifies what can be expected from constraints alone and where genuinely new patterns emerge in networked systems changing continuously over time.
Data and Code Availability
All code used to generate the figures and reproduce the experiments is publicly available on github/Maximum-Entropy-Temporal-Networks and Zenodo. All figures are fully reproducible from the repository.